Determine the mean and variance of the random variable with the following probability mass function f (x) (27/13)(1/3)*, x 1,2,3 Round your answers to three decimal places (e.g. 98.765) Mean- Variance
Answer #1

mean E(X)=\sumxP(x)=1*(27/13)*(1/3)+2*(27/13)*(1/3)2+3*(27/13)*(1/3)3 =1.385

variance =\sum(x-mean)2*P(x)=(1-1.385)2*(27/13)*(1/3)+(2-1.385)2*(27/13)*(1/3)2+(3-1.385)2*(27/13)*(1/3)3 =0.391 ( please try 0.390 if this comes wrong due to rounding)